OpenStudy (anonymous):

PLEASE HELP Simplify [(9x^2-1)/(8x-4)] / [(3x^2+5x-2)/(2x^2+1x-1)]

OpenStudy (anonymous):

so far I've simplified it to \[(3x^2)/(4x-4) \times (x^2+1x-1)/(x^2+5x-2) \]

OpenStudy (anonymous):

\[\large{\frac{9x^2-1}{8x-4} \div \frac{3x^2+5x-2}{2x^2+x-1}}\] Factor each trinomial: \[\large{\frac{(3x+1)(3x-1)}{4(2x-1)} \div \frac{3x^2+5x-2}{2x^2+x-1}}\] That's a start. I'll be back if you want assistance.

OpenStudy (anonymous):

But don't you have to change is to a multiplication problem by multiplying the first equation by the inverse of the second?

OpenStudy (anonymous):

Yes. You could invert the second fraction before factoring.

OpenStudy (anonymous):

okay, thanks!

OpenStudy (anonymous):

I'll post my answer after you post yours.... :-)

OpenStudy (anonymous):

okay, wait I'm stuck because I don't know what to do with (3x+1)(3x-1) do I foil it or..?

OpenStudy (anonymous):

No, once you have factored all parts, then try to cancel anything that you can.

OpenStudy (anonymous):

okay so those 3x's cancel out with the 3x from 3x^2+5x-2?

OpenStudy (anonymous):

No, first you have to factor.

OpenStudy (anonymous):

Just like I did for the first fraction, you have to factor the second. Do you know how to factor a trinomial?